OpenStudy (anonymous):

Simplify: 3√5 + √5 Can somebody simplify this, and explain to me how to do a problem like this? Thanks.

OpenStudy (anonymous):

do you know how to collect like terms?

OpenStudy (anonymous):

Yes. Why?

OpenStudy (anonymous):

\[\sqrt{5}\] is a like term

OpenStudy (anonymous):

let x = sqrt 5 3x + x = ....

OpenStudy (anonymous):

I wasn't sure, but that's what I was going to do. Are you saying the answer is 3√10?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

look above ^

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

if 3x + x = 4x and x = sqrt 5 then 4sqrt5 is your ans

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Explain a bit?

OpenStudy (anonymous):

3sqrt5 + 1sqrt5 is basically 3 x sqrt5 + 1 x sqrt5 so that means (3 + 1) x sqrt5 so that equals to 4sqrt5

OpenStudy (anonymous):

the x 's are multiplication signs

OpenStudy (anonymous):

Oh. Damn, that's simple. Can't believe I didn't understand it before. Well thanks.

OpenStudy (anonymous):

np